OpenStudy (anonymous):

The tickets for a high school basketball game are $2.50 for students and $4.00 for adults. At Tuesday night's game, 95 tickets were sold for a total of $290.00. How many student tickets were sold?

OpenStudy (anonymous):

2.50*x+(95-x)*4=290

OpenStudy (anonymous):

How to solve algebraically: s = number of student tickets a = number of adult tickets s + a = 95 2.5s + 4a = 290 Solve the system.

OpenStudy (anonymous):

here the value of students ticket is assumed as x

OpenStudy (anonymous):

a = 95 -s 2.5s + 4(95 -s) = 290 2.5s + 380 - 4s = 290 -1.5s = -90 s = 60 60 tickets. It doesn't matter which variables you use, I just used s and a to make it easier to remember what each variable means.

OpenStudy (anonymous):

so the last part to get 60 you divided?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

If the value of x is assumed, then the value of the students tickets are to be assumed as x..... Therefore, let a + s=95..... and 2.5s + 4a= 290. Solving using the systems of equations, s=95-a... a=95-s, then, using substitution, 2.5s +4(95-s)=290, becomes, s=60... thus, there are 60 tickets.

OpenStudy (anonymous):

Yeah, @cyclevleoce12345 explains it better but that's the gist of it. You create a system from the given information and then solve it.

OpenStudy (anonymous):

thankyou guys

OpenStudy (anonymous):

np, happy to help (: